### v2

# Detection of new very-high-energy sources outside the galactic plane in the Fermi-LAT data

Abstract
We present a search for spatio-temporal clusters in 16 years of Fermi-LAT very-high-energy (VHE; $E>100$ GeV) data using the DBSCAN algorithm, focusing on high Galactic latitude ($|b|>10^{\circ}$) clusters with $\geq 5$ events and transient doublets (two events within $\leq 3$ days). Of 107 detected clusters, two correspond to previously unidentified VHE sources: weak BL Lacertae objects 4FGL J0039.1-2219 and 4FGL J0212.2-0219, promising targets for next-generation VHE observatories.
Due to the low VHE photon background, even doublets with a duration of several days exhibited high statistical significance. While most of the 114 detected doublets originated from bright TeV emitters (e.g., Mrk 421, Mrk 501), we identified six VHE flares lacking TeVCat associations. Five of these flares correlate with sources from the Third Catalog of Fermi-LAT High-Energy Sources (3FHL), while one ’orphan’ flare lacks a high-energy (HE; $E>10$ GeV) source counterpart. Some of these flares reached extreme luminosities of $\mathcal{O}(10^{47}~\mathrm{erg~s^{-1}})$.
No consistent temporal correlation emerged between HE and VHE activity: HE flares preceded, coincided with, or followed VHE emission across sources, with some showing no HE counterpart. Remarkably, 3FHL J0308.4+0408 (NGC 1218) is a Seyfert Type I galaxy, while no object of this class was known as a VHE emitter before. The ’orphan’ flare without any known HE source in the vicinity may originate from NGC 5549, a low-luminosity LINER galaxy. Both sources expand the limited sample of non-blazar AGN detected at VHE energies. The fact that some weak sources with non-aligned jets and, sometimes, even without any traces of HE activity, could demonstrate very short and powerful VHE flares cannot be easily accounted in many current AGN models and calls for their further development.
keywords:
very-high-energy gamma-rays, blazars, BL Lacs, AGNs high-energy transients
journal: High Energy Astrophysics
[…]
[1]organization=Sternberg Astronomical Institute, Moscow M.V. Lomonosov State University, addressline=Universitetskiy Prospekt 13, city=Moscow, postcode=119991, country=Russia
[…]
[2]organization=Lebedev Physical Institute, Astro Space Center, Pushchino Radio Astronomy Observatory, addressline=Radioteleskopnaya 1a, city=Pushchino, Moscow reg., postcode=142290, country=Russia
[…]
[3]organization=The Institute for Nuclear Research of the Russian Academy of Sciences, addressline=B-312, Prospekt 60-letiya Oktyabrya, 7a, city=Moscow, postcode=117312, country=Russia
[…]
[4]organization=Faculty of Physics, Moscow M.V. Lomonosov State University, addressline=Leninskie gory 1, city=Moscow, postcode=119991, country=Russia
1 Introduction
Despite significant observational progress in recent decades, the very-high-energy (VHE; $E>100$ GeV) sky remains understudied compared to other wavelengths and energy ranges. Imaging atmospheric Cherenkov telescopes (IACTs) – the most sensitive instruments at these energies – unfortunately have very small fields of view (FoV), which makes conducting deep all-sky surveys extremely challenging. Observations by the Fermi Large Area Telescope (Fermi-LAT; Atwood et al., 2009) could be used to partially alleviate the problem: although its effective area is relatively modest ($\mathcal{O}(1~\mathrm{m^{2}})$), Fermi-LAT achieves a substantial total exposure – approximately $\mathcal{O}(10^{12}~\mathrm{cm^{2}~s})$ – owing to its exceptionally wide field of view (2.4 sr) and uninterrupted operation over 16 years. For context, it is equivalent to a few hours of HESS observations, given that at these energies, HESS has an effective area of $\mathcal{O}(10^{4}~\mathrm{m^{2}})$ (Aharonian et al., 2006). Furthermore, Fermi-LAT’s high cadence of observations provides a unique capability to detect short-lived VHE flares, including rare ’orphan’ flares that lack counterparts at other electromagnetic wavelengths. This combination of long time span of observation, all-sky coverage, and high temporal resolution makes Fermi-LAT a critical tool for studying transient and variable VHE phenomena.
In this paper, we present a search for VHE $\gamma$-ray clusters and transients at Galactic latitudes $|b|>10^{\circ}$ using 16 years of Fermi-LAT data, focusing on out-of-plane sources likely of extragalactic origin. Many of these sources belong to the high-synchrotron-peak (HSP) blazar class (Tavecchio et al., 1998), where particles accelerated to extreme energies in active galactic nucleus (AGN) jets produce a spectral energy distribution (SED) dominated by synchrotron emission peaking in X-rays and inverse-Compton scattering at VHE energies. These extreme astrophysical accelerators are not only critical for AGN studies (Costamante et al., 2001) but also serve as probes for the intergalactic magnetic fields (Tavecchio et al., 2011; Khalikov and Dzhatdoev, 2021), the extragalactic background light (H. E. S. S. Collaboration et al., 2017), and physics beyond the Standard Model (Abdalla et al., 2021).
Fermi-LAT’s continuous monitoring capabilities have been leveraged for VHE source detection since the mission’s early phases: an analysis of just 1.5 years of data yielded a significant detection of VHE emission from the galaxy IC 310 (Neronov et al., 2010) and 5 new Galactic sources at $|b|<10^{\circ}$ (Neronov and Semikoz, 2010). In 5 years of observations, Fermi-LAT detected thirteen high-redshift ($z>0.5$) sources (Neronov et al., 2015) and 2 VHE photons from the distant ($z=1.1$) blazar PKS 0426-380 (Tanaka et al., 2013). Subsequent studies expanded searches to lower energies (tens of GeV or below) using clustering algorithms such as the minimal spanning tree (MST; Campana et al., 2008) and density-based spatial clustering (DBSCAN; Tramacere and Vecchio, 2013). These efforts, combined with accumulating data, produced the 12-year MST catalog of 1888 $\gamma$-ray source candidates above 10 GeV at $|b|>20^{\circ}$ (Campana, R. et al., 2018; Campana and Massaro, 2021), identified eight new candidates of supernova remnants in the Large Magellanic Cloud (Tramacere et al., 2025), and revealed 18 sources without counterparts in the Fermi-LAT sources catalog (Soor et al., 2025).
Despite these advances, the wealth of Fermi-LAT VHE data remains underexplored. Here, we apply the DBSCAN algorithm to systematically search for transients and clustered $\gamma$-ray signals, prioritizing candidates unassociated with known VHE sources. This approach aims to uncover new phenomena and refine existing constraints on extreme astrophysical environments.
The paper is organized as follows. In Section 2 we describe our selection of data and the algorithms that we used in our paper. In Section 3 we present our results and we conclude in Section 4.
2 Data selection and algorithm of cluster detection
2.1 Fermi-LAT dataset
We analyzed 16 years of Fermi-LAT data spanning 4 August 2008 to 9 August 2024, selecting SOURCE-class events with energies >100>100 GeV. The initial dataset contained 47,764 photons. To mitigate contamination from the Galactic diffuse gamma-ray background, we excluded events within $|b|<10^{\circ}$ of the Galactic plane, as this region exhibits significantly higher background flux that reduces sensitivity to faint sources. After applying this latitude cut, the final sample comprised 21,686 VHE photons.
2.2 Description of the DBSCAN method
DBSCAN is a density-based clustering method that uses the local density of data points to find clusters in data sets taking into account the background noise. The algorithm groups data points, utilizing two key parameters: $\varepsilon$ (the maximum distance defining the neighborhood of a point) and $K$ (the minimum number of points required to form a cluster core). For each point, the algorithm counts the number of neighbors within its $\varepsilon$-neighborhood. If a point satisfies the core condition (i.e. the number of neighbors is greater than or equal to $K$, including itself), it is designated as a core point and initiates a new cluster including the neighboring points. Clusters are formed by connecting core points and their reachable neighbors. Points within the $\varepsilon$-neighborhood of a core point but with fewer than $K$ neighbors themselves are called border points. They are part of a cluster but cannot expand it. Points unassociated with any cluster are classified as noise. We implemented DBSCAN via the scikit-learn Python package [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.DBSCAN.
As a first approximation we choose $K$ and $\varepsilon$ as follows:
  1. 1.
    Set $\varepsilon=0.3^{\circ}$ which is close to the instrumental point spread function (PSF) at a given threshold energy (for $E>100$ GeV, $\theta_{\mathrm{PSF}}=0.1^{\circ}$)
  2. 2.
    To minimize false positives, we derived $K$ using Poisson statistics. For the average background level $\lambda$:
    $\lambda=N/\Omega_{|b|>10^{\circ}},$ (1)
    where $N=21686$ is the total number of analyzed photons, $\Omega_{|b|>10^{\circ}}=4\pi(1-\sin{10^{\circ}})$ is the solid angle of the survey, the probability that $k$ photons fall within a solid angle $\Omega_{\varepsilon}$ is described by the Poisson distribution:
    $P_{\mathrm{loc}}(k)=\dfrac{(\lambda\Omega_{\varepsilon})^{k}\exp{(-\lambda\Omega_{\varepsilon})}}{k!}$ (2)
    Finally, we used the Bonferroni correction, accounting for a large number of trials, so the global false-alarm probability became:
    $P_{\mathrm{glob}}(k)=\dfrac{\Omega_{|b|>10^{\circ}}}{\Omega_{\varepsilon}}P_{\mathrm{loc}}(k)$ (3)
    where $\Omega_{|b|>10^{\circ}}/\Omega_{\varepsilon}$ was the number of analyzed fields. For the given $\lambda$ and $\varepsilon$ $P_{\mathrm{glob}}(k)$ depends on $k$ as follows:
    $k$ 4 5 6 7
    $P_{\mathrm{glob}}$ 4.39 0.158 0.00474 0.00012
    Based on this, we adopted $K=5$ ($P_{\mathrm{glob}}\approx 0.16$) for an initial clustering. Final candidate clusters were re-evaluated individually for their statistical significance.
  3. 3.
    The average background level was calculated for each cluster using the following procedure: we considered a $10^{\circ}\times 10^{\circ}$ square centered on the center of the cluster $(l_{0},~b_{0})$ (if $|b_{0}|<20^{\circ}$, we take $b\in[10^{\circ},~20^{\circ}]$ or $b\in[-20^{\circ},~-10^{\circ}]$, if $b_{0}<0^{\circ}$), after that we excluded all photons within $0.5^{\circ}$ of any 3FHL source in this square and obtain a new number of background photons $N^{\prime}$ and an average background level $\lambda^{\prime}$.
    The refined background density $\lambda^{\prime}$ and the cluster solid angle $\Omega_{\mathrm{cls}}=\pi\rho^{2}_{\mathrm{max}}$ (where $\rho_{\mathrm{max}}$ is the cluster’s maximum angular radius) were substituted into Eq. 3 to calculate updated $P_{\mathrm{glob}}$ values.
2.3 Extended cluster analysis
Clusters exhibiting significant spatial extension may originate from non-pointlike sources. To quantify this, we compare their photon distributions to the distribution of photons from the Crab Nebula and pulsar (close to a point source for our purposes at very high energies, with a radius of $\sim 1.5^{\prime}$ (Aharonian et al., 2024)). The Crab reference cluster contains 376 photons (210 front-converted, 166 back-converted), providing an empirical template for the Fermi-LAT point-spread function (PSF) at $E>100$ GeV.
  1. 1.
    Photon Properties and Empirical PSF: For each photon in a cluster, we calculate its angular distance $r$ to the candidate source position (or cluster centroid). The probability $p_{i}(r)$ of observing a photon of conversion type $i$ (front/back) at a distance $\geq r_{0}$ is: $p_{i}(r)=N_{i}(r\geq r_{0})/N_{i}^{\mathrm{total}}$,
    where $N_{i}(r\geq r_{0})$ is the number of type-$i$ photons in the Crab cluster beyond $r_{0}$, and $N_{i}^{\mathrm{total}}$ is the total number of the type-$i$ photons in the Crab cluster.
  2. 2.
    Likelihood Analysis: For a test cluster with $N_{0}$ front and $N_{1}$ back photons, we compute the log-likelihood value:
    $\mathcal{L}=-\left[\sum_{j=1}^{N_{0}}\ln{p_{0}(r_{j})}+\sum_{k=1}^{N_{1}}\ln{p_{1}(r_{k})}\right]$ (4)
    To assess significance, we generate a likelihood distribution by randomly drawing 10,000 mock samples of $N_{0}$ and $N_{1}$ photons from the Crab cluster and calculating $\mathcal{L}$ for each.
  3. 3.
    Extension Significance: The p-value for a point source is: $P_{\mathrm{ps}}=N(\mathcal{L}>\mathcal{L}_{0})/N_{\mathrm{tot}}$, where $\mathcal{L}_{0}$ is the likehood value for the analyzed cluster, $N(\mathcal{L}>\mathcal{L}_{0})$ is the number of mock samples from the Crab distribution with likelihoods exceeding $\mathcal{L}_{0}$, and $N_{\mathrm{tot}}=10,000$ is the total number of mock samples. A low $P_{\mathrm{ps}}<0.01$ suggests significant extension.
2.4 Transients
While Section 2.2 focused on spatial clustering, we now incorporate temporal constraints by requiring photon arrival times to lie within a 3-day window. This modification accounts for the transient nature of potential sources. We chose this temporal scale in order to select short and bright VHE flares, which have a typical duration of days.
For temporal clustering, the background rate $\lambda^{*}$ is adjusted to reflect the reduced time window:
$\lambda^{*}=\lambda^{{}^{\prime}}\frac{\delta t_{\mathrm{win}}}{T_{\mathrm{obs}}}$ (5)
where $\delta t_{\mathrm{win}}=3$ days and $T_{\mathrm{obs}}\approx 16$ years is the total observation time. This scaling gives the expected background count within a 3-day interval. Using Equation (3) with $\lambda^{*}$, the revised threshold for cluster detection becomes $K=2$ (a ’doublet’), as even two temporally coincident photons are statistically significant at this low level of background. The significance ($P_{\mathrm{glob}}$) of each candidate transient is recalculated using the methodology outlined in Section 2.2, now incorporating both spatial and temporal proximity.
3 Results
3.1 Clusters
Our analysis identified 107 clusters with a global significance threshold of $P_{\mathrm{glob}}<0.05$. All of these clusters have an associated source from the Fourth Fermi Large Area Telescope catalog (4FGL; Abdollahi et al., 2020). To prioritize novel VHE emitters, we imposed a stringent angular separation criterion, retaining only clusters located $1^{\circ}$ from any source in the Third Catalog of Fermi-LAT High-Energy Sources (3FHL; Ajello et al., 2017) or TeVCat [^2] https://www.tevcat.org/ (Wakely and Horan, 2008) (online interactive catalog for VHE sources).This produced three candidate VHE sources (Table 2) with no prior associations from VHE catalogs (3FHL and TeVCat respectively).
Most clusters showed no evidence of spatial extension, with photon distributions consistent with the Fermi-LAT PSF ($P_{\mathrm{ps}}>0.01$). The sole exception is Cluster # 91 ($N_{\mathrm{ph}}=9,~P_{\mathrm{glob}}=2.54\times 10^{-4}$, centered at $l=126.92^{\circ},~b=13.02^{\circ}$ and associated with 4FGL J0153.0+7517 (Fig. 1)). This cluster exhibits marginal extension, with $P_{\mathrm{ps}}=6.8\times 10^{-3}$ when referenced to its centroid and $P_{\mathrm{ps}}^{\mathrm{3FHL}}=0.033$ when anchored to the nearest 3FHL source position.
Notably, our method successfully recovered known extended sources, including the supernova remnant CTA 1 and the Large Magellanic Cloud (LMC), validating its sensitivity to resolved structures.
Figure 1: Spatial distribution of $\gamma$-ray photons associated with the source 4FGL J0153.0+7517. The cluster’s centroid is indicated by a red cross symbol, corresponding to Galactic coordinates $l=126.92^{\circ},~b=13.02^{\circ}$
3.1.1 4FGL J0039.1-2219
This source is a low-redshift ($z=0.064$), low-luminosity BL Lacertae (BL Lac) blazar included in the 3HSP catalog (Chang et al., 2019) – a comprehensive compilation of very high-energy (VHE) candidate sources selected using a multi-frequency dataset. The observed VHE flux is consistent with the power-law index and flux extrapolated from lower-energy measurements reported in 4FGL. The angular distribution of $\gamma$-ray events around the source is shown in Fig. 2.
Figure 2: Spatial distribution of $\gamma$-ray photons associated with the source 4FGL J0039.1-2219. The cluster’s centroid is indicated by a red cross symbol, corresponding to Galactic coordinates $l=93.36^{\circ},~b=-84.35^{\circ}$
3.1.2 4FGL J1544.3-0649
4FGL J1544.3-0649 (Fig. 3) is a blazar candidate first detected following a major $\gamma$-ray flare in May 2017 (Bruni et al., 2018). The source may represent a rare transient blazar with episodic high-energy (HE) and VHE emission and an exceptionally low duty cycle (Sahakyan and Giommi, 2021). Alternatively, it could be an AGN exhibiting intermittent emission from secondary mini-jets whose orientations are misaligned relative to the primary jet axis (Shao et al., 2022). It is evident that this source passed our cuts only because of its transient nature, which prevented its entering into the 3FHL, TeVCat and 3HSP catalogs as they were compiled prior to its flare.
Figure 3: Spatial distribution of $\gamma$-ray photons associated with the source 4FGL J1544.3-0649. The cluster’s centroid is indicated by a red cross symbol, corresponding to Galactic coordinates $l=0.30^{\circ},~b=36.15^{\circ}$
3.1.3 4FGL J0212.2-0219
This blazar (Fig. 4), located at $z=0.25$, is cataloged in the 3HSP (Chang et al., 2019) and was previously analyzed as a candidate for Imaging Atmospheric Cherenkov Telescopes (IACTs) in (Nievas Rosillo et al., 2022). While deemed too faint for detection by next-generation ground-based telescopes (e.g., CTA), our analysis reveals a marginal excess of very-high-energy (VHE) photons compared to predictions based on the spectral parameters reported in the 4FGL catalog. This discrepancy suggests the source’s VHE flux may deviate from the extrapolated power-law spectrum, leaving open the possibility of detectability under improved sensitivity or during flaring episodes.
Figure 4: Spatial distribution of $\gamma$-ray photons associated with the source 4FGL J0212.2-0219. The cluster’s centroid is indicated by a red cross symbol, corresponding to Galactic coordinates $l=164.34^{\circ},~b=-58.59^{\circ}$
3.2 Transients
Blazars exhibit significant variability across their electromagnetic spectra, also in the VHE range. While Fermi-LAT data have been used to study VHE variability in bright sources (Paliya, 2024), we focus on extreme, short-term variability (flares) in weaker sources where traditional light-curve analysis is impractical due to low photon counts. Such flares could manifest as spatially and temporally clustered VHE photons, where even doublets (two photons within a few days) can achieve high statistical significance given the low background at VHE. We identified 114 doublets of events separated by not more than 3 days (259200 seconds) and 0.3 degrees. Most of them originate from bright sources, such as Mrk 421 (57 doublets), Mrk 501 (16), 1ES 1959+650 (11), BL Lac (5), 1ES 1218+304 (4), PKS 2155-304 (4), PG 1553+113 (3), 3FHL J0508.0+6737333A designation in the 3FHL catalog is given when there is no TeVCat association (2), 1ES 1011+496 (1), 1ES 1727+502 (1), 3FHL J1457.8-4642 (1), 3FHL J0308.4+0408 (1), 3FHL J0847.0-2337 (1), 3FHL J1203.3-3924 (1), 3FHL J1037.6+5711 (1), 3C 66A (1), PKS 1440-389 (1), PKS 0447-439 (1), S5 0716+714 (1), and one orphan flare without counterparts in the 4FGL, 3FHL and TeVCat catalogs. Additionally, we detected 12 triplets (10 from Mrk 421, 2 from 1ES 1959+650) and 1 quadruplet (Mrk 421). We focused on the 6 doublets from sources without TeVCat associations or prior VHE flare reports (excluding 3FHL J0508.0+6737, which exhibited a VERITAS-detected flare in 2009 (Ong, 2009)). These doublets show exceptional significance ($P_{\mathrm{glob}}\in(10^{-8},~10^{-4})$), even without accounting for the spatial coincidence with 3FHL sources (Table 3).
Given the low statistics, two photons constrain the flare’s expected event count to $0.35<N_{flare}<6.5$ (95% CL). This translates to the VHE flux limits: $0.17F_{VHE}^{0}<F_{VHE}<3.25F_{VHE}^{0}$, where $F_{VHE}^{0}=2/\mathcal{E}$ and $\mathcal{E}$ is the Fermi-LAT exposure during the doublet’s time window. For each candidate flare, we analyzed all VHE photons within $0.3^{\circ}$ of the associated source and and 10-100 GeV photons [^4] We will refer to them as HE photons in the following. in this circle in the time interval ($t_{0}-10^{7}$ s, $t_{0}+10^{7}$ s), where $t_{0}$ is the doublet’s midpoint.
3.2.1 3FHL J0308.4+0408 (NGC 1218)
The VHE and HE photon lists for this source are provided in Appendix Tables 4 and 5. Only two VHE photons were detected within $0.3^{\circ}$ of the source over 16 years, implying a flare with a flux increase of $10^{3}$ relative to the quiescent VHE level. Notably, only one of 36 HE photons coincided with the flare epoch (observed 5.5 Ms prior to the VHE event), suggesting an orphan VHE flare without a HE counterpart.
NGC 1218 is a S0 galaxy hosting the radio-bright Seyfert Type I nucleus 3C 78. While Seyfert galaxies sometimes exhibit GeV flaring activity (Gokus et al., 2021), VHE flares have been reported only in the Type II Seyfert NGC 1275 (Perseus Cluster core; (MAGIC Collaboration et al., 2018; Cao et al., 2024)). Intriguingly, recent IceCube observations associate Seyferts like NGC 1068 and NGC 7469 with astrophysical neutrinos (IceCube Collaboration et al., 2022; Sommani et al., 2025), raising questions about their potential as transient multi-messenger sources.
3.2.2 3FHL J0847.0-2337 (PMN J0847-2337)
PMN J0847-2337 is a candidate blazar exhibiting an unusual temporal pattern: a GeV-flare detected after a VHE one. Analysis of its photon distribution (Tables 6, 7) reveals that 8 of 90 HE photons arrived within a 3 Ms window starting 4 Ms after the VHE flare, with 5 photons concentrated in a 10-day sub-interval. Naturally, there is a strong observational bias against detection of such flares with IACTs, which usually exploit HE flares as a sort of external triggers.
3.2.3 3FHL J1037.6+5711 (GB6J1037+5711)
This relatively bright BL Lac object exhibits 10 VHE and 325 HE photons in total. The VHE flare coincides with a prolonged period of enhanced HE activity (Tables 8, 9), and a transient event was independently detected by HAWC approximately one week after the Fermi-LAT doublet (Weisgarber and HAWC Collaboration, 2018).
Despite its brightness, this source lacks a secure redshift measurement (Table 3). Spectroscopic observations in SDSS DR7 reported $z=0.83\pm 0.0004$ but with low confidence ($z_{\mathrm{conf}}=0.07$) (Abazajian et al., 2009). Follow-up observations with larger telescopes revealed a featureless optical spectrum, yielding only a lower limit of $z>0.25$ (Paiano et al., 2020; Kasai et al., 2023). Photometric estimates from broadband SED modeling gave a less reliable but intriguing value of $z=1.14$ (Foschini et al., 2022).
The source’s AGN-dominated emission ($\sim 0.5$ mag brightening over two decades) further obscures host galaxy features, complicating the redshift determination.
3.2.4 3FHL J1203.3-3924 (PMN J1203-3926)
This source is a blazar candidate, possibly of the BL Lac type. It exhibited a VHE flare without a HE counterpart, and none of the 40 HE photons arrived in the 20 Ms interval of interest (see Table 10).
3.2.5 3FHL J1457.8-4642 (PMN J1457-4642)
Only two VHE photons were detected from this blazar candidate during the entire 16-year time span of the observations, both constituting a temporally clustered doublet. This VHE activity was preceded by enhanced HE emission: 4 out of the 12 HE photons arrived within a 1.5 Ms window starting 3 Ms prior to the VHE event (Tables 11, 12).
This HE precursor to VHE flaring is a common feature of blazar variability. For instance, during the 2011 outburst of the BL Lacertae, HE activity preceded VHE emission by $\sim$ month (Arlen et al., 2013).
3.2.6 NGC 5549?
The lack of a 3FHL/4FGL counterpart complicates the identification of the host source for this VHE doublet. While several galaxies reside within the same galaxy group as NGC 5549 – a low-luminosity AGN exhibiting faint nuclear activity – NGC 5549 itself emerges as the most plausible candidate due to its weak but discernible AGN signatures.
The $0.2^{\circ}$ region around the doublet’s centroid contains only distant quasars ($z>0.75$), which are unlikely counterparts given their extreme redshifts and spectral characteristics. Strikingly, not only VHE flare (see Table 13) lacks any HE counterpart: zero HE photons were detected in $0.3^{\circ}$ in all 16 years of observations.
3.2.7 Energetics and spectral properties of the flares
The VHE luminosities of the flaring sources exhibit extraordinary variability, increasing by up to three orders of magnitude compared to their quiescent states. During peak activity, some flares reached VHE(!) luminosities of $\mathcal{O}(10^{47}~\mathrm{erg~s^{-1}})$, rivaling the bolometric output of luminous quasars. For context, we estimate the VHE luminosities of two well-studied blazars during their brightest 3-day intervals – Mrk 421, $1.7\times 10^{46}~\mathrm{erg~s^{-1}}$ (MET 240723290–240962846; 18–20 August 2008, 4 photons), and BL Lacertae: $6.0\times 10^{45}~\mathrm{erg~s^{-1}}$ (MET 647619845 – 647750055; 10– 12 July 2021, 2 photons). These values are an order of magnitude lower than the most extreme flares identified in this work, underscoring the exceptional nature of the newly detected events. Several flares lacked detectable HE counterparts, implying extremely hard spectra. This is corroborated by the energies of VHE photons in key doublets: 3FHL J1203.3-3924 (172 GeV and 203 GeV photons), NGC 5549 (orphan flare; 156 GeV and 593 GeV photons).
An interesting representative of this class could be the VHE quadruplet from Mrk 421 (4 photons within 3 days), which occurred without contemporaneous HE or X-ray activity (Abdo et al., 2011; Bartoli et al., 2016). This event presents two hypotheses, either there was a missed HE precursor – a hypothetical HE flare $\sim 1$ month prior could have escaped detection due to Fermi-LAT’s commissioning phase at the time, or it was just a statistical fluctuation. Given that in 5840 days there were observed 372 VHE photons from this source, the probability of a random 4-photon cluster with a 180 ks duration is $9\%$ (Poissonian estimate).
The total absence of HE emission could pose a challenge to the standard synchrotron self-Compton (SSC) model, suggesting that some alternative mechanisms of emission could be responsible for such flares.
Detected flares bear striking resemblance to several instances of fast variability of another misaligned sources, i.e. IC 310 and M 87 (Rieger and Levinson, 2018; Rieger, 2019): in both cases we need some new additional VHE component to account for the observed spectrum. Still, if we took quiescent HE luminosity as a proxy for a jet power $L_{j}$ (Nemmen et al., 2012), VHE -jet luminosity ratios $L_{\mathrm{VHE}}/L_{j}$ have considerably higher values for our set, which possibly disfavor models with magnetospheric origin of the emission and various jet-object interaction scenarios.
4 Summary and conclusions
In this paper we searched for spatio-temporal clusters of the VHE events ($E>100$ GeV) in 16 years of the Fermi-LAT data at galactic latitudes $|b|>10^{\circ}$. Using the DBSCAN algorithm, we found 107 clusters. All, but three of them, were associated with already known VHE sources, listed in the 3FHL or TeVCat catalogs (Table LABEL:tab:clusters_full_list). From remaining three (Table 2), 4FGL J1544.3-0649, which had an epoch of activity in 2017-2018, could belong to a rare class of transient blazars and because of that evaded inclusion into VHE catalogs. The other sources, 4FGL J0039.1-2219 and 4FGL J0039.1-2219 are weak BL Lacs, that could be a tip of an iceberg of much wider population of weaker sources, that would be detected with the next-generation IACTs.
We also searched for flaring VHE activity, looking for event multiplets within short temporal intervals of 3 days. We found 114 doublets, most of them (57) from Mrk 421 – the brightest extragalactic source with a very high level of variability. We focused on the 6 doublets without TeVCat sources in their proximity. Five of them were associated with 3FHL sources (J0308.4+0408, J0847.0-2337, J1037.6+5711, J1203.3-3924, J1457.8-4642), and one has no association with any known HE/VHE sources (see Table 3).
Due to the limited effective area of the Fermi-LAT this approach is sensitive only to the brightest flares with a very high luminosity that in cases of 3FHL J1037.6+5711 and 3FHL J1457.8-4642 could reach $\mathcal{O}(10^{47}~\mathrm{erg~s^{-1}})$ (isotropic), which is, e.g., an order of magnitude larger than the brightest BL Lac flare in July 2021.
One of the sources – 3FHL J0308.4+0408(NGC 1218) – is a Seyfert Type I galaxy. No objects of this type not belonging to narrow-line Sy 1 subclass were known as VHE sources before. A fortiori, no short energetic VHE flares were anticipated from such objects. Even more striking is the plausible association of the ’orphan’ doublet with the LINER galaxy NGC 5549. This type of galaxies are considered to harbor low-luminosity AGNs in their centers and currently they are below sensitivity limit of HE/VHE telescopes. Still, there was a major extremely hard flare, during which the VHE flux ($E>100$ GeV) exceeded the average flux at HE ($E>10$ GeV) at least a thousandfold, and the VHE luminosity of the AGN surpassed $10^{45}~\mathrm{erg~s^{-1}}$. It is worth noting, that non-blazar AGNs are a small minority in the full population of TeV sources, with only 4 Fanaroff-Riley I galaxies present in the TeVCat (M87, 3C 264, NGC 1275, and Cen A), so a detection of VHE flares from low-luminous non-aligned AGNs could have considerable significance for the field of AGN jet studies.
HE behaviour during these flares was quite diverse: 3FHL J1037.6+5711 had coinciding VHE/HE flares, the HE flare in 3FHL J1457.8-4642 preceded the VHE one by about a month, there was HE activity 1.5 month after the VHE flare in 3FHL J0847.0-2337, which is quite uncommon, and in 3FHL J0308.4+0408, 3FHL J1203.3-3924, NGC 5549? there were no HE counterparts whatsoever.
Note: When this paper was in a final stage of preparation, another work in which a DBSCAN algorithm was used to search for clusters in GeV Fermi data, appeared (Cozzolongo et al., 2025). Their analysis focused on identifying extended sources in the Galactic plane, contrasting with our emphasis on extragalactic transients and high-latitude sources.
Acknowledgements
The study was funded by a grant from the Russian Science Foundation 25-12-00111 (https://rscf.ru/en/project/25-12-00111/). This research has made use of NASA’s Astrophysics Data System and the SIMBAD database, operated at CDS, Strasbourg, France.
Appendix A Table of clusters
# $l$, $b$, $N_{\text{ph}}$ $P_{\text{glob}}$ Nearest 4FGL $\rho_{\text{4FGL}},\ ^{\circ}$ Nearest 3FHL $\rho_{\text{3FHL}},\ ^{\circ}$
0 179.84 65.04 372 J1104.4+3812 0.00715 J1104.4+3812 0.00496
1 76.96 21.77 10 $3.46\times 10^{-11}$ J1838.8+4802 0.05751 J1838.8+4802 0.05598
2 165.06 -31.75 7 $6.35\times 10^{-06}$ J0319.8+1845 0.07170 J0319.8+1845 0.07020
3 17.71 -52.27 87 $2.45\times 10^{-166}$ J2158.8-3013 0.02860 J2158.8-3013 0.02726
4 102.90 -24.76 6 $3.25\times 10^{-05}$ J2322.7+3436 0.03764 J2322.6+3436 0.03232
5 98.32 -18.50 7 $2.78\times 10^{-05}$ J2250.0+3825 0.08845 J2250.0+3825 0.08653
6 298.75 -15.82 6 $3.49\times 10^{-03}$ J1130.5-7801 0.09545 J1130.5-7801 0.09011
7 150.23 -13.73 6 $1.28\times 10^{-05}$ J0316.8+4120 0.05194 J0316.6+4120 0.05965
8 18.22 -14.33 14 $1.11\times 10^{-17}$ J1917.7-1921 0.02275 J1917.7-1921 0.01938
9 29.48 68.20 35 $1.16\times 10^{-65}$ J1427.0+2348 0.00959 J1427.0+2348 0.01108
10 63.59 38.85 124 $2.24\times 10^{-259}$ J1653.8+3945 0.00947 J1653.8+3945 0.00883
11 143.79 15.94 42 $3.83\times 10^{-44}$ J0507.9+6737 0.04702 J0508.0+6737 0.04749
12 195.36 -19.61 7 $2.19\times 10^{-08}$ J0509.4+0542 0.04642 J0509.4+0542 0.04134
13 191.78 -33.18 8 $3.89\times 10^{-08}$ J0416.9+0105 0.04268 J0416.8+0105 0.02835
14 282.74 -40.76 5 $9.84\times 10^{-05}$ J0353.0-6831 0.04569 J0353.0-6832 0.05355
15 248.79 -39.93 43 $3.59\times 10^{-85}$ J0449.4-4350 0.01883 J0449.4-4350 0.02069
16 21.91 44.00 99 $1.64\times 10^{-167}$ J1555.7+$11$ 0.03366 J1555.7+$11$ 0.03280
17555This clusters contains 2 close sources: J1221.3+3010 (44 photons) and J1217.9+3006 (17 photons) 187.06 82.53 61 $5.93\times 10^{-124}$ J1221.3+3010 0.23119 J1221.3+3010 0.22455
18 77.35 64.88 11 $1.40\times 10^{-17}$ J1428.5+4240 0.06910 J1428.5+4240 0.06110
19 98.03 17.67 152 $1.15\times 10^{-236}$ J2000.0+6508 0.02509 J1959.9+6508 0.02636
20 306.04 19.78 5 $4.71\times 10^{-02}$ J1307.6-4259 0.03644 J1307.6-4259 0.03435
21 165.54 52.72 36 $1.29\times 10^{-63}$ J1015.0+4926 0.01263 J1015.0+4926 0.01314
22 3.26 10.59 7 $9.76\times 10^{-04}$ J1714.0-2029 0.06513 J1714.0-2028 0.06933
23 249.67 18.49 12 $6.73\times 10^{-12}$ J0912.9-2102 0.09428 J0912.9-2103 0.08537
24 350.39 -32.59 23 $9.21\times 10^{-27}$ J2009.4-4849 0.01434 J2009.4-4849 0.00996
25 190.27 10.96 36 $2.00\times 10^{-57}$ J0650.7+2503 0.03895 J0650.7+2503 0.03922
26 56.78 11.92 5 $8.36\times 10^{-04}$ J1850.5+2631 0.02421 J1850.4+2631 0.04292
27 166.27 32.91 19 $4.81\times 10^{-37}$ J0809.8+5218 0.02201 J0809.8+5218 0.01780
28 11.02 -26.72 6 $2.53\times 10^{-03}$ J1958.3-3010 0.08546 J1958.3-3011 0.09158
29 247.77 12.19 11 $8.34\times 10^{-14}$ J0847.0-2336 0.01326 J0847.0-2337 0.01002
30 131.91 45.63 13 $1.97\times 10^{-26}$ J1136.4+7009 0.01486 J1136.5+7009 0.01732
31 320.77 -11.15 5 $3.40\times 10^{-05}$ J1610.7-6648 0.02619 J1610.6-6649 0.01945
32 278.33 -60.78 12 $1.25\times 10^{-17}$ J0209.3-5228 0.01143 J0209.3-5229 0.01831
33 225.46 67.36 8 $1.39\times 10^{-11}$ J1117.0+2013 0.05154 J1117.0+2014 0.04253
34 144.03 28.04 24 $7.76\times 10^{-36}$ J0721.9+7120 0.04577 J0721.8+7120 0.04924
35 93.96 -81.35 5 $2.93\times 10^{-06}$ J0033.5-1921 0.13761 J0033.5-1921 0.13991
36 140.17 -16.79 30 $4.37\times 10^{-36}$ J0222.6+4302 0.02852 J0222.6+4302 0.02812
37 112.89 -10.07 6 $2.09\times 10^{-02}$ J2347.0+5141 0.15451 J2347.0+5142 0.15805
38 214.61 -60.20 10 $1.48\times 10^{-20}$ J0303.4-2407 0.02414 J0303.4-2407 0.02177
39 309.50 19.42 9 $6.77\times 10^{-07}$ J1325.5-4300 0.03100 J1325.5-4300 0.02866
40 243.45 -19.99 8 $3.16\times 10^{-07}$ J0627.0-3529 0.02001 J0627.1-3528 0.02425
41 60.26 63.68 5 $2.65\times 10^{-09}$ J1448.0+3608 0.03231 J1447.9+3608 0.04511
42 261.35 -11.34 5 $4.08\times 10^{-07}$ J0746.6-4754 0.03600 J0746.6-4755 0.02292
43 132.46 -22.96 8 $1.38\times 10^{-07}$ J0136.5+3906 0.03699 J0136.5+3906 0.03962
44 266.93 20.05 6 $1.73\times 10^{-10}$ J1010.2-3119 0.01853 J1010.2-3119 0.02442
45 157.46 25.40 5 $2.58\times 10^{-04}$ J0710.4+5908 0.05572 J0710.4+5908 0.06219
46 350.09 -67.56 6 $5.71\times 10^{-12}$ J2324.7-4041 0.04388 J2324.7-$40$ 0.04433
47 351.14 -26.99 5 $2.88\times 10^{-04}$ J1936.9-4720 0.03668 J1936.9-4720 0.02770
48 240.67 -43.58 5 $2.33\times 10^{-04}$ J0428.6-3756 0.04557 J0428.6-3756 0.04820
49 152.42 -18.41 6 $7.11\times 10^{-08}$ J0312.9+3614 0.00644 J0312.8+3614 0.01734
50 138.92 30.81 18 $3.39\times 10^{-30}$ J0805.4+7534 0.02880 J0805.5+7534 0.02682
51 263.52 -31.48 24 $1.25\times 10^{-36}$ J0543.9-5531 0.00615 J0543.9-5532 0.00509
52 123.75 58.79 6 $2.56\times 10^{-09}$ J1248.3+5820 0.02066 J1248.3+5820 0.01891
53 278.42 -53.12 8 $2.44\times 10^{-09}$ J0244.6-5819 0.03614 J0244.4-5819 0.03664
54 150.58 -13.28 15 $2.72\times 10^{-21}$ J0319.8+4130 0.01751 J0319.8+4130 0.01879
55 47.45 -30.22 5 $4.84\times 10^{-02}$ J2104.3-0212 0.11744 J2104.2-0212 0.11968
56 107.05 -61.87 5 $5.24\times 10^{-05}$ J0022.0+0006 0.07363 J0022.0+0006 0.05967
57 307.51 20.06 7 $4.08\times 10^{-06}$ J1315.0-4236 0.04673 J1315.0-4237 0.05338
58 161.42 54.42 9 $3.11\times 10^{-15}$ J1031.3+5053 0.03815 J1031.3+5053 0.03205
59 67.76 -10.79 6 $4.35\times 10^{-05}$ J2042.1+2427 0.02653 J2042.0+2428 0.03159
60 167.84 66.11 15 $1.07\times 10^{-22}$ J1120.8+4212 0.05440 J1120.8+4212 0.05593
61 279.56 -31.72 22 $2.00\times 10^{-22}$ J0537.8-6909 0.02272 J0537.9-6909 0.01731
62 119.73 10.81 8 $8.05\times 10^{-03}$ J0007.0+7303 0.35210 J0007.0+7303 0.34733
63 13.09 -78.09 5 $2.09\times 10^{-07}$ J2359.0-3038 0.09019 J2359.1-3038 0.09073
64 283.89 74.51 7 $3.37\times 10^{-13}$ J1230.8+1223 0.02985 J1230.8+1223 0.03775
65 347.88 24.23 8 $4.08\times 10^{-05}$ J1548.8-2250 0.06035 J1548.7-2250 0.06277
66 340.69 27.48 9 $1.89\times 10^{-06}$ J1517.7-2422 0.09283 J1517.6-2422 0.10043
67 201.92 -45.70 5 $1.31\times 10^{-06}$ J0349.4-1159 0.01386 J0349.3-1159 0.01245
68 153.96 14.23 5 $4.49\times 10^{-03}$ J0540.5+5823 0.03701 J0540.5+5823 0.04052
69 325.66 18.72 20 $1.30\times 10^{-39}$ J1443.9-3908 0.02015 J1443.9-3908 0.01497
70 145.70 43.10 5 $5.60\times 10^{-05}$ J0958.7+6534 0.04408 J0958.7+6533 0.04688
71 133.51 80.54 5 $3.91\times 10^{-07}$ J1243.2+3627 0.09365 J1243.2+3627 0.08915
72 92.60 -10.44 44 $1.41\times 10^{-66}$ J2202.7+4216 0.00625 J2202.7+4216 0.01138
73 233.80 -17.58 7 $1.07\times 10^{-11}$ J0622.3-2605 0.02166 J0622.4-2606 0.03592
74 100.77 -13.23 7 $1.49\times 10^{-08}$ J2247.8+4413 0.04598 J2247.9+4413 0.04050
75 201.89 18.13 5 $8.21\times 10^{-06}$ J0738.1+1742 0.06056 J0738.1+1742 0.06771
76 351.25 40.15 7 $7.07\times 10^{-05}$ J1512.8-0906 0.03401 J1512.8-0906 0.02758
77 34.19 24.50 11 $3.01\times 10^{-10}$ J1725.0+1152 0.06263 J1725.0+1152 0.06015
78 77.11 33.55 14 $1.70\times 10^{-20}$ J1728.3+5013 0.03547 J1728.3+5013 0.04306
79 119.75 17.80 7 $5.76\times 10^{-06}$ J2340.8+8015 0.08647 J2340.8+8015 0.08861
80 18.62 -21.00 8 $1.43\times 10^{-05}$ J1944.9-2143 0.10907 J1944.9-2143 0.10566
81 67.32 13.40 7 $2.33\times 10^{-04}$ J1904.1+3627 0.03917 J1904.1+3627 0.04055
82 229.58 -66.30 6 $7.19\times 10^{-08}$ J0238.4-3116 0.06258 J0238.4-3117 0.06161
83 87.38 73.46 6 $2.32\times 10^{-07}$ J1341.2+3958 0.06589 J1341.2+3959 0.05339
84 166.05 27.27 5 $3.50\times 10^{-03}$ J0733.4+5152 0.04548 J0733.4+5152 0.04788
85 327.77 14.64 8 $7.36\times 10^{-06}$ J1503.6-4146 0.02214 J1451.8-4145 2.18606
86 282.17 -61.32 5 $2.12\times 10^{-04}$ J0156.9-5301 0.06521 J0156.7-5302 0.07746
87 293.36 -17.56 5 $3.28\times 10^{-06}$ J0953.4-7659 0.05297 J0953.3-7659 0.04906
88 33.37 70.59 6 $6.57\times 10^{-08}$ J1417.9+2543 0.10932 J1418.0+2543 0.11528
89 213.44 -83.44 6 $1.49\times 10^{-12}$ J0120.4-2701 0.09762 J0120.4-2701 0.10540
90 325.24 19.34 7 $6.09\times 10^{-09}$ J1440.6-3846 0.04781 J1440.6-3846 0.04649
91 126.92 13.02 9 $2.54\times 10^{-04}$ J0153.0+7517 0.13220 J0152.8+7517 0.13704
92 134.12 39.25 6 $1.08\times 10^{-04}$ J1031.1+7442 0.08638 J1031.2+7442 0.07792
93 93.36 -84.35 5 $7.91\times 10^{-07}$ J0039.1-2219 0.15912 J0033.5-1921 3.13907
94 95.90 -52.36 6 $1.23\times 10^{-07}$ J2346.7+0705 0.06576 J2346.6+0705 0.05724
95 276.86 -32.66 5 $4.08\times 10^{-02}$ J0531.8-6639e 0.16534 J0531.8-6639e 0.16534
96 100.56 30.83 5 $1.57\times 10^{-03}$ J1748.6+7005 0.12346 J1748.6+7006 0.12108
97 358.32 -48.23 6 $3.34\times 10^{-02}$ J2139.4-4235 0.09523 J2139.4-4234 0.10106
98 151.77 51.78 10 $1.19\times 10^{-18}$ J1037.7+5711 0.00843 J1037.6+5711 0.01053
99 107.43 -26.10 5 $2.70\times 10^{-02}$ J2343.6+3438 0.08011 J2343.6+3439 0.06839
100 129.94 -37.22 6 $7.57\times 10^{-05}$ J0115.8+2519 0.06736 J0115.8+2519 0.06616
101 0.30 36.15 7 $2.48\times 10^{-12}$ J1544.3-0649 0.02103 J1549.9-0659 1.41385
102 164.34 -58.59 5 $2.98\times 10^{-06}$ J0212.2-0219 0.05389 J0217.8+0143 4.33239
103 323.26 -34.37 5 $9.16\times 10^{-03}$ J2040.2-7115 0.16504 J2040.3-7116 0.17108
104 93.31 19.74 6 $6.45\times 10^{-07}$ J1926.8+6154 0.00103 J1926.9+6154 0.00839
105 152.90 -36.74 6 $1.57\times 10^{-05}$ J0232.8+2018 0.16848 J0232.8+2017 0.16124
106 8.58 59.99 5 $3.29\times 10^{-04}$ J1442.7+1200 0.19075 J1442.8+1200 0.20314
Appendix B Table of VHE cluster candidates
# $l,^{\circ}$ $b,^{\circ}$ $N_{\mathrm{ph}}$ $P_{\mathrm{glob}}$ 4FGL name $F_{1-100}$ $\gamma$ $z$
93 93.36 -84.35 5 $7.9\times 10^{-7}$ J0039.1-2219 2.1 1.8 0.064
101 0.30 36.15 7 $2.5\times 10^{-12}$ J1544.3-0649 12.1 1.9 0.17
102 164.34 -58.59 5 $3\times 10^{-6}$ J0212.2-0219 3.1 2.0 0.25
Table 2: Properties of unassociated VHE cluster candidates and their associations in 4FGL. Columns: (1) Cluster ID from Table LABEL:tab:clusters_full_list; (2,3) Galactic longitude and latitude of cluster centroid; (4) Photon count; (5) Global probability of chance occurrence; (6) Associated 4FGL source; (7) Integrated GeV flux from 4FGL in $10^{-10}~\mathrm{cm^{2}~s^{-1}}$; (8) Spectral index; (9) Redshift.
Appendix C Table of transients
[…]
3FHL name $\rho$, Assoc. name $z$ type $F_{1-100}$ $P_{\text{glob}}$ $\Delta t$ $\mathcal{E}$ $F_{VHE}^{0}$ $L$
J0308.4+0408 0.013 NGC 1218 0.03 Sy 1 0.83 $8.0\times 10^{-9}$ 1.82 4.4 4.5 1.8
J0847.0-2337 0.12 PMN J0847-2337 0.059 BLL 1.27 $1.3\times 10^{-5}$ 1.65 2.7 7.4 11.6
J1037.6+5711 0.05 GB6J1037+5711 ?1 BLL 4.08 $1.7\times 10^{-6}$ 1.72 5.7 3.5 ?
J1203.3-3924 0.10 PMN J1203-3926 0.28 BLL 0.44 $5.5\times 10^{-5}$ 1.38 1.4 14.2 503
J1457.8-4642 0.10 PMN J1457-4642 0.16 BLL? 0.17 $1.2\times 10^{-4}$ 1.49 1.6 12.5 145
0.007 NGC 5549? 0.025 LIN $3.3\times 10^{-4}$ 2.16 3.2 6.2 1.7
  • 1
    there are several redshift estimation for GB6J1037+5711, we discuss it in more details in the pertinent section below
Appendix D Lists of events for transients
$t$, s $E$, GeV $l$, $b$, Distance,
560582016 115.9 174.840 -44.514 0.013
560764691 121.9 174.841 -44.511 0.012
Table 4: List of VHE photons for 3FHL J0308.4+0408 (NGC 1218). Bold rows indicate the doublet events. Columns: (1) Mission elapsed time (seconds); (2) Energy (GeV); (3,4) Galactic coordinates; (5) Angular distance from source center (degrees).
$t$, s $E$, GeV $l$, $b$, Distance,
555018647 54.0 174.857 -44.554 0.04
Table 5: List of 10-100 GeV photons for 3FHL J0308.4+0408 (NGC 1218) in $2\times 10^{7}$ s interval centered on doublet. Columns are the same as in Table 4 Events, temporally coinciding with the doublet, are shown in boldface.
$t$, s $E$, GeV $l$, $b$, Distance,
257661702 491.4 247.803 12.184 0.03
363076249 106.8 247.665 12.145 0.12
370057437 135.9 247.703 12.195 0.07
448525521 291.9 247.926 12.149 0.16
478319482 207.6 247.738 12.270 0.08
541749219 189.0 247.771 12.254 0.05
575995109 103.7 247.789 12.205 0.01
655392252 204.6 247.615 12.252 0.16
655557713 130.2 247.731 12.268 0.08
689089028 136.0 247.911 11.979 0.26
728372459 156.4 247.815 12.217 0.04
Table 6: List of VHE photons for 3FHL J0847.0-2337 (PMN J0847-2337), analogous to Table 4
$t$, s $E$, GeV $l$, $b$, Distance,
648643223 20.8 247.661 12.191 0.11
659048631 88.6 247.906 12.286 0.15
659088756 17.1 247.829 11.993 0.22
659316050 26.9 247.787 12.157 0.05
659636820 11.0 247.635 12.318 0.18
659881515 93.8 247.742 12.269 0.07
661263355 14.0 247.816 12.237 0.05
662222468 45.2 247.743 12.229 0.04
662302983 13.8 247.564 12.177 0.21
Table 7: List of 10-100 GeV photons for 3FHL J0847.0-2337 (PMN J0847-2337) analogous to Table 5.
$t$, s $E$, GeV $l$, $b$, Distance,
486429674 126.6 151.735 51.806 0.03
511813676 133.3 151.767 51.813 0.03
515627358 135.0 151.816 51.823 0.04
518976121 118.8 151.714 51.817 0.05
551111293 126.1 151.708 51.773 0.04
551283348 169.7 151.696 51.823 0.06
584157729 118.9 151.800 51.763 0.03
650447843 104.0 151.747 51.736 0.05
654983499 133.0 151.822 51.822 0.05
735040791 245.4 151.888 51.652 0.15
Table 8: List of VHE photons for 3FHL J1037.6+5711 (GB6J1037+5711), analogous to Table 4
$t$, s $E$, GeV $l$, $b$, Distance,
542669658 23.3 151.758 51.834 0.05
546965261 25.0 151.717 51.802 0.04
546998880 30.4 151.662 51.851 0.10
547089708 51.1 151.808 51.806 0.03
547552717 17.3 151.844 51.778 0.04
549223585 64.3 151.809 51.799 0.03
549798866 11.0 151.459 51.705 0.21
550164818 17.4 151.820 51.788 0.03
551152678 11.3 151.379 51.886 0.26
551401970 62.1 151.682 51.696 0.10
551607468 10.9 151.745 51.754 0.03
553354233 10.2 151.479 51.964 0.26
553993869 16.3 151.726 51.796 0.03
554927761 27.2 151.758 51.838 0.06
555345377 16.8 151.833 51.724 0.07
555517519 13.7 151.679 51.767 0.06
557514043 12.3 151.670 51.756 0.07
558648352 10.8 151.640 51.691 0.12
558744923 37.3 152.147 51.617 0.29
Table 9: List of 10-100 GeV photons for 3FHL J1037.6+5711 (GB6J1037+5711) analogous to Table 5.
$t$, s $E$, GeV $l$, $b$, Distance,
260094483 101.2 293.114 22.563 0.21
373271798 172.2 292.963 22.533 0.07
373409583 203.3 292.754 22.634 0.19
694676662 101.9 292.877 22.501 0.02
Table 10: List of VHE photons for 3FHL J1203.3-3924 (PMN J1203-3926), analogous to Table 4
$t$, s $E$, GeV $l$, $b$, Distance,
391732356 158.7 324.526 10.815 0.15
391881102 148.2 324.357 11.010 0.16
Table 11: List of VHE photons for 3FHL J1457.8-4642 (PMN J1457-4642), analogous to Table 4
$t$, s $E$, GeV $l$, $b$, Distance,
388126965 17.7 324.369 10.825 0.02
388309964 55.8 324.365 10.840 0.01
389170862 30.9 324.383 10.801 0.05
389618335 10.8 324.459 10.743 0.13
Table 12: List of 10-100 GeV photons for 3FHL J1457.8-4642 (PMN J1457-4642) analogous to Table 5.
$t$, s $E$, GeV $l$, $b$, Distance,
580411826 156.0 353.234 61.364 0.04
580628550 592.9 353.25 61.437 0.03
Table 13: List of VHE photons for NGC5549(?), analogous to Table 4